Suppose U and V are subspace of R^n. Prove that orthogonal of ( U intersection V)=orthogonal of U+ orthogonal of V

1
Expert's answer
2021-12-26T16:30:32-0500

for "u,v\isin U\cap V" :

if "x\isin (U\cap V)^{\perp}" and "U\cap V \neq 0" then:

"x\cdot u=0" or "x\cdot v=0" , so

"x\isin U^{\perp}+V^{\perp}"

so,

"(U\cap V)^{\perp}" is subset of "U^{\perp}+V^{\perp}"


if "x\isin U^{\perp}+V^{\perp}" then:

"x\cdot u=0" or "x\cdot v=0" for "u\isin U" and "v\isin V"

then "x\isin (U\cap V)^{\perp}"

so,

"U^{\perp}+V^{\perp}" is subset of "(U\cap V)^{\perp}"


that is, "(U\cap V)^{\perp} =U^{\perp}+V^{\perp}"